OpenStudy (anonymous):

4-x^2>or=0 How do I graph this on a number line?

OpenStudy (anonymous):

first you have to solve the inequality. you have \[4-x^2\geq 0\] \[(2-x)(2+x)\geq 0\] so the solution is \[-2\leq x\leq 2\]

OpenStudy (anonymous):



OpenStudy (anonymous):

I'm not sure how factoring helps.\[4-x^{2}\ge0\]\[x^{2}\le4\]\[-2 \le x \le 2\]